OpenStudy (kimo_indie):

What is the equation of a line parallel to the line y = 4x + 7 and passes through the point (3, 0)

OpenStudy (anonymous):

As the line is parallel, the gradient (m) is the same. m = 4 (y - b) = m (x - a) (y - 0) = 4 (x - 3) y = 4x - 12